OpenStudy (anonymous):

Find the area between two curves y=e^2x and y=sin2x where 0 < x< (pi/2) - -

OpenStudy (danjs):

is that e^(2x) or e^2 * x and sin^2(x) or sin(2x)

OpenStudy (anonymous):

e^(2x) and sin(2x)

OpenStudy (anonymous):

The one problem is seem to be having is finding out when e^(2x)-sin(2x)=0 The rest I can do.

OpenStudy (anonymous):

Hello?

jimthompson5910 (jim_thompson5910):

you'll need to use a graphing calculator to find the approximate solutions to e^(2x)-sin(2x)=0

jimthompson5910 (jim_thompson5910):

I don't think it's possible to solve e^(2x)-sin(2x)=0 without using a calculator

OpenStudy (anonymous):

something wrong here

OpenStudy (anonymous):

they intersect a bunch of places to the left of zero

OpenStudy (anonymous):

The calculator is giving me the wrong area, its going over regions that are not suppose to be shaded.

OpenStudy (anonymous):

also the answer has to be within 0 and pi/2

OpenStudy (anonymous):

sine is periodic, and \(e^{2x}\) has a horizontal asymptote at the x axis, so they intersect infinitely often

OpenStudy (anonymous):

which is why I was confused on what to do since there is no real answer for e^2x-sin2x=0. Because if I get the answer to that then I know how to set up the second integral

jimthompson5910 (jim_thompson5910):

you are looking for this green area (see attached image) the red curve is the function y = sin(2x) the blue curve is y = e^(2x) the bounds on x are 0 < x < pi/2

jimthompson5910 (jim_thompson5910):

so you just need to compute the integrals from 0 to pi/2 (for the two functions) and subtract

OpenStudy (anonymous):

Yes, when I do that on my TI-84 the area beneath the curve in between x=0 and x=1 is being shaded as well

OpenStudy (anonymous):

I keep getting the answer to 10. something when its suppose to be 8. something

jimthompson5910 (jim_thompson5910):

it's not between x = 0 and x = 1 it's between x = 0 and x = pi/2

OpenStudy (anonymous):

okay thanks

OpenStudy (danjs):

half e to the pi minus 3 halves

OpenStudy (danjs):

\[\frac{ e^\pi }{ 2 }-\frac{ 3 }{ 2 }\]

OpenStudy (anonymous):

how do I show my work for this?

OpenStudy (danjs):

Calculate 2 definite integrals for x from zero to half pi

OpenStudy (anonymous):

Yeah I understand that one goes from pi/2 to 0. And the other goes from pi/2 to what number?

OpenStudy (danjs):

\[\int\limits_{0}^{\pi/2}e^{2x}dx - \int\limits_{0}^{\pi/2}\sin(2x)dx\]

OpenStudy (anonymous):

that 's what I'm having trouble with, in other to find that other number I have to find the answer to this equation: e^2x-sin2x=0. But since there is no answer to this , that why I'm confused

OpenStudy (anonymous):

are you sure thats how you do it? because this is why my professor told us to do

OpenStudy (danjs):

the two functions don't intersect on that interval

jimthompson5910 (jim_thompson5910):

you only do e^2x-sin2x=0 when there are points of intersection (between x = 0 and x = pi/2)

jimthompson5910 (jim_thompson5910):

so you know which function is above the other (and where this occurs)

OpenStudy (anonymous):

\[\int\limits_{0}^{\pi/2}e^2x-(\sin(2x)\]

OpenStudy (anonymous):

i thought thats how I'm suppose to do this problem

OpenStudy (anonymous):

and the second equation you would add to the previous one

jimthompson5910 (jim_thompson5910):

you have it correct, you can break it up like dan is showing

OpenStudy (anonymous):

\[\int\limits_{\pi/2}^{don't know this}\]

jimthompson5910 (jim_thompson5910):

it's from 0 to pi/2, so you do know the upper piece

jimthompson5910 (jim_thompson5910):

the graph shows no intersections between x = 0 and x = pi/2

OpenStudy (anonymous):

but the answer that I am getting 10.1 is not right the answer should be 8. something

OpenStudy (danjs):

\[\frac{ e^{2x} }{ 2 }-\frac{- \cos(2x) }{ 2 }\] that is what the integrals evaluate to, now just use upper bound pi/2 and lower bound zero, to calculate the value for the definate integral

jimthompson5910 (jim_thompson5910):

I'm getting 10.07035 roughly not sure how it's 8 something

OpenStudy (anonymous):

yeah that's what I got too but my professor said the answer is 8 something

jimthompson5910 (jim_thompson5910):

something seems missing if that's the case

jimthompson5910 (jim_thompson5910):

not sure what

OpenStudy (anonymous):

yeah that's why I thought that there should be a seccond equation

jimthompson5910 (jim_thompson5910):

can you post a screenshot of the full question?

OpenStudy (anonymous):

here ill show you an example from my online homework

OpenStudy (anonymous):

yeah I will

OpenStudy (danjs):

You get 8.167 if you integrate from x=1 to x=pi/2 and add them instead of subtract. lol i dont know what you looking for.

OpenStudy (anonymous):

im so sorry, for the confusion. But here is the question

OpenStudy (anonymous):

its problem 8b

jimthompson5910 (jim_thompson5910):

yeah the answer is 10.07035 (approximately) there is either a typo on the sheet or the professor is mistaken

OpenStudy (anonymous):

oh okay! Thank for so much for your help! I appreciate it greatly. And sorry for not understanding that quickly.

OpenStudy (danjs):

yeah. The first one 8a uses that process you described, setting the two functions equal to find the bounds for integration, -2 to 3

jimthompson5910 (jim_thompson5910):

that's ok, you picked it up just fine you're welcome